kenzie126:

4X -(-2x – 6) = 6(x + 1)

Shadow:

Is this just solving for x?

kenzie126:

yeah

jhonyy9:

so just distribute these parentheses and calcule the x value

kenzie126:

i forgot how to do that

Shadow:

\[a(b + c) = ab + ac\]

justus:

4x+2x+6=6(x+1) 4x+2x+6=6x+6 4x-4x+6=6 -4x=-4x Now divide each term by -4 and simplify to get your final answer

kenzie126:

so im dividing 6 by 4

justus:

No.

kenzie126:

what am i dividing ?

justus:

the last numbers -4x and 4x

mxddi3:

4x+2x+6=6x+6 going back to this, i thought we would have 4x+2x=6x so on the left we have 6x+6 and on the right we had 6x+6 then you simplify by combining like terms.... im not sure wym by tht Justus.

mxddi3:

\(\color{#0cbb34}{\text{Originally Posted by}}\) justus 4x+2x+6=6(x+1) 4x+2x+6=6x+6 4x-4x+6=6 -4x=-4x Now divide each term by -4 and simplify to get your final answer \(\color{#0cbb34}{\text{End of Quote}}\) ok they would get the same answer i think, nvm

justus:

Yeah they do, it's fine. :)

justus:

So now -4x and 4x, divide them by -4